OpenStudy (anonymous):

Checking account A charges a monthly service fee of $2 and a per-check fee of $0.30, while checking account B charges a monthly service fee of $3 and a per-check fee of $0.20. How many checks would a person have to write for the two accounts to cost the same? i really need help with this!

OpenStudy (anonymous):

The way to work this problem is to start by setting up expressions for how much each account costs. For example, account A charges $2 no matter how many checks, plus 0.30 for each check. You could write this like: COST = 2 + 0.30x where x is the number of checks.

OpenStudy (anonymous):

Can you get the expression for account B? COST = ____ + ______x

OpenStudy (anonymous):

ok i think i got what your saying

OpenStudy (anonymous):

Ok :) In words, the cost of account A is equal to the flat fee of $2 plus the number of checks multiplied by 30 cents per check. The cost of account B is a larger flat fee of $3 plus the number of checks multiplied by a smaller per check amount of only 20 cents.

OpenStudy (anonymous):

Say you only write 1 check... Account A costs $2 + ($0.30 * 1 check) = $2.30 Account B costs $3 + ($0.20 * 1 check) = $3.20 What if you write 2 checks? Account A costs $2 + ($0.30 * 2 checks) = $2.60 Account B costs $3 + ($0.20 * 2 checks) = $3.40

OpenStudy (anonymous):

The way to find out the number of checks for when the costs of each account is the same is to set those two expressions equal. Cost of Account A = Cost of Account B 2 + 0.30x = 3 + 0.20x

OpenStudy (anonymous):

thank you so much that really helped(:

OpenStudy (anonymous):

Do you see how to go from there? You can get the x's on one side and the numbers on the other, then solve for x. That "x" is the number of checks that makes the account costs the same. When you get x, make sure to check your work by getting the account costs for A and for B and make sure they really are the same :)